# Impact of the continuum Coulomb interaction in quantum-orbit-based treatments of high-order above-threshold ionization

### 08 Dec 2023 v1

@@ IV.2 Photoelectron momentum distributions
In Fig. 8 we present the photoelectron momentum distributions for the shortest [(panel (a)] and the second-shortest [(panel (b)] forward-scattering solutions of the SP equations. The upper half of the panels corresponds to the SFA distributions, while the lower-half panels are calculated using the CQSFA theory. The results shown in different panels are normalized independently. For the shortest pair of solutions, the photoelectron momentum distributions calculated using our theories qualitatively agree well. However, the positions of the minima caused by the interference of two contributions are not at the same place. This is particularly the case for the minima with a small energy. The fact that the interference minima do not appear at the same place in the photoelectron momentum plane can be explained in the similar manner as for the backward-scattering solutions. In addition, for the emission close to the polarization direction, the CQSFA calculation leads to an increase in the photoelectron yield in comparison with the one obtained using the SFA theory. Finally, similarly to the backward-scattering solutions, the CQSFA does not give any nonzero probability density after the cutoff because the complex trajectories are not included.
@@
In Fig. 8(b) we see that, for the second-shortest pair of solutions, the differences between the presented results are much more pronounced than for the shortest pair. This is expected due to the fact that the contribution of this solution is significant only in the low-energy part of the spectra. In conclusion, due to the fact that the contributions of the forward-scattering solutions are significant for low- or medium-energy regions, the Coulomb effects should be taken into consideration.
@@ V Conclusions
In this work, we compare two semianalytical theories based on the saddle-point method which allow one to get a clear insight into the underlying physics of the strong-field ionization process. The first one is the strong-field approximation (SFA) which assumes that the driving field is so strong that the Coulomb attraction between the liberated electron and the residual ion can be neglected during the electron propagation in the continuum. The second theory is the Coulomb quantum-orbit strong-field approximation (CQSFA) for which the Coulomb potential and the driving-field potential are treated equally. These theories are structurally very different. The former is a Born-type-expansion theory for which the rescattering times and conditions are well-defined because the interactions with the parent ion during rescattering are localized at the single point, namely the origin. On the other hand, the latter has a Coulomb distorted continuum which consequently makes it impossible to clearly define when the rescattering starts or ends.
@@
To compare the results obtained using the SFA and CQSFA is not as trivial task as it sounds at first. In particular, for a systematic comparison, it is not sufficient to remove the Coulomb coupling or make it short-range due to the structural differences between the two theories. These differences make it hard to establish a one-to-one correspondence between the orbits obtained with each method. In order to make this comparison, we have identified the CQSFA trajectories which mimic the behavior of the Born-type SFA trajectories with a single act of rescattering. This means that among all orbits which appear in the CQSFA theory, we have extracted only those with SFA counterparts.
@@
We have found that the agreement between the results obtained using our theories is fairly good if the solutions used in the CQSFA theory are restricted to only those which we expect to have a counterpart in the SFA theory. This happens because the compared orbits accumulate similar phases in the continuum, although their dynamics can be very different. For example, before the rescattering, the SFA trajectories are one-dimensional for a linearly polarized driving field, while for the CQSFA theory, the corresponding orbits are always two-dimensional. More generally, the SFA trajectories follow the field, which is not the case with their CQSFA counterparts due to the always present Coulomb interaction.

### 21 Feb 2024 v2

@@ IV.2 Photoelectron momentum distributions
In Fig. 8 we present the photoelectron momentum distributions for the shortest [(panel (a)] and the second-shortest [(panel (b)] forward-scattering solutions of the SP equations. The upper half of the panels corresponds to the SFA distributions, while the lower-half panels are calculated using the CQSFA theory. The results shown in different panels are normalized independently. For the shortest pair of solutions, the photoelectron momentum distributions calculated using our theories qualitatively agree well. However, the positions of the minima caused by the interference of two contributions are not at the same place. This is particularly the case for the minima with a small energy. The fact that the interference minima do not appear at the same place in the photoelectron momentum plane can be explained in a similar manner as for the backward-scattering solutions. In addition, for the emission close to the polarization direction, the CQSFA calculation leads to an increase in the photoelectron yield in comparison with the one obtained using the SFA theory. Finally, similarly to the backward-scattering solutions, the CQSFA does not give any nonzero probability density after the cutoff because the complex trajectories are not included.
@@
In Fig. 8(b) we see that, for the second-shortest pair of solutions, the differences between the presented results are much more pronounced than for the shortest pair. This is expected due to the fact that the contribution of this solution is significant only in the low-energy part of the spectra. In conclusion, since the contributions of the forward-scattering solutions are significant for low- or medium-energy regions, the Coulomb effects should be taken into consideration.
IV.3 Probing a wider parameter range
In this section we check if and how the derived conclusions about the SFA and CQSFA results depend on the driving-field parameters, for the first and second shortest pair of orbits. First, we investigate how the difference between the SFA and CQSFA saddle-point solutions for the ionization and rescattering times depends on the laser-field parameters.
In Fig. 9 we present the ionization [panels (a) and (c)] and rescattering [panels (b) and (d)] times for the driving-field intensity W/cm and wavelengths of 2000 nm [pink (bright gray) lines], 1300 nm [orange (light gray) lines], and 800 nm [red (gray) lines]. In addition, the blue (dark gray) lines correspond to the saddle-point solutions obtained using the driving field with intensity W/cm and wavelength of 800 nm. The dashed (solid) lines denote the solutions obtained using the SFA (CQSFA) theory. Figures 9(a) and (b) [Figures 9(c) and (d)] correspond to the shortest (second-shortest) pair of the SP solutions. The change of the driving-field parameters affects the difference between the saddle-point solutions obtained using our two theories which is particularly visible for the ionization time. In particular, as the driving-field wavelength increases, the difference between the results becomes less pronounced. For example, the difference between the ionization-time SP solutions for the wavelength of 800 nm is far larger than for the wavelength of 2000 nm, for the same laser-field intensity [cf. the difference between the red (gray) solid and dashed lines and the pink (bright gray) solid and dashed lines in Figs. 9(a) and (c)]. This is more pronounced for the shortest pair of the SP solutions. Moreover, the difference between the SFA and CQSFA results is less affected by the change in the driving field intensity [cf. the difference between the red (gray) solid and dashed lines and the blue (dark gray) solid and dashed lines in Figs. 9(a) and (c)]. This happens because the electron excursion in the laser field is directly proportional to the square root of the driving-field intensity and directly proportional to the square of the applied-field wavelength. Because the Coulomb effects fade away at the large distance of the core, the agreement between the SFA and CQSFA results becomes better much more rapidly with the increase of the wavelength than with the increase of the intensity. Nonetheless, there are subtle differences observed for different driving-field intensities, in particular for the second shortest pair of orbits. For the lower intensity, the range of ionization times is slightly narrower, and the low-energy ends of the arch-like structures associated with the rescattering times close at marginally larger values. This is due to the increased influence of the Coulomb potential. Finally, we note that the curves which correspond to the ionization and rescattering times as functions of the photoelectron energy are closed regardless of the values of the driving-field parameters for the second-shortest pair of the SP solutions. As the photoelectron energy approaches zero, the two solutions coalesce. This coalescence is not sudden since the photoelectron requires a certain minimum energy to escape the Coulomb interaction. The reason for this behaviour being restricted to the second-shortest pair (it also occurs for the fourth-shortest pair which is not shown in Fig. 9) was elaborated on earlier using the results shown in Fig. 1(c).
Let us now investigate how the difference between the photoelectron momentum distributions obtained using the SFA and CQSFA theories is affected by the change of the applied-field parameters. In Fig. 10 we present the photoelectron momentum distributions for the shortest [panels (a) and (c)] and the second-shortest [panels (b) and (d)] solutions of the SP equations calculated using the SFA (upper-half of the panels) and the CQSFA (lower-half of the panels) theories. The driving-field intensity and wavelength are W/cm and nm [panels (a) and (b)], and W/cm and nm [panels (c) and (d)], respectively. For both cases, the agreement between the calculated results is good, and the differences are mainly restricted to narrow regions in the photoelectron momentum plane. However, the agreement is better for the larger intensity and longer wavelength [cf. panel (a) with panel (c), and panel (b) with panel (d)]. For the case of the long wavelength and the high intensity, the minima caused by the interference of the contributions of the two solutions of one pair coincide very well for both the shortest and second-shortest pair. The only difference appears in the narrow region around for the shortest and around for the second-shortest pair. On the other hand, for the smaller intensity and shorter wavelength, the agreement is not excellent, particularly for the second-shortest pair of the SP solutions in which case the minima obtained using the SFA and CQSFA theories do not appear at the same place in the photoelectron momentum plane. Finally, we mention that we did not choose the laser field with the wavelength of 2000 nm for our illustration since in this case, the number of interference minima is so large that it is difficult to access the agreement between the results obtained using our theories.
@@ V Conclusions
In this work, we compare two semianalytical theories based on the saddle-point method, which allows one to get a clear insight into the underlying physics of the strong-field ionization process. The first one is the strong-field approximation (SFA) which assumes that the driving field is so strong that the Coulomb attraction between the liberated electron and the residual ion can be neglected during the electron propagation in the continuum. The second theory is the Coulomb quantum-orbit strong-field approximation (CQSFA) for which the Coulomb potential and the driving-field potential are treated equally. These theories are structurally very different. The former is a Born-type-expansion theory for which the rescattering times and conditions are well-defined because the interactions with the parent ion during rescattering are localized at a single point, namely the origin. On the other hand, the latter has a Coulomb distorted continuum which consequently makes it impossible to clearly define when the rescattering starts or ends.
@@
To compare the results obtained using the SFA and CQSFA is not as trivial a task as it sounds at first. In particular, for a systematic comparison, it is not sufficient to remove the Coulomb coupling or make it short-range due to the structural differences between the two theories. These differences make it hard to establish a one-to-one correspondence between the orbits obtained with each method. In order to make this comparison, we have identified the CQSFA trajectories that mimic the behavior of the Born-type SFA trajectories with a single act of rescattering. This means that among all orbits that appear in the CQSFA theory, we have extracted only those with SFA counterparts. We also provide an example of an orbit for which this correspondence does not hold, although CQSFA orbits with no SFA counterparts have been studied in far more detail in previous publications [54, 55, 59, 60].
Here, we have also focused on the simplest possible system: hydrogen in a linearly polarized monochromatic field. This was purposefully done not to introduce other effects, which would mask what we intended to single out. Focal averaging may distort the interference patterns of interest [64, 82], or be a source of incoherence [62]. Short pulses will exhibit unequal cycles, which would give rise to rescattering ridges of unequal energies [61, 83] and holographic patterns that may look quite different from those obtained with longer pulses or monochromatic fields [84, 85, 86]. Averaging over the carrier-envelope phase will also weaken or wash out subtler interference patterns [87, 88]. Bi- or polychromatic fields would break symmetries and lead to additional solutions to those in Table 1, even for very weak fields [79]. These distortions increase for two-color fields with comparable amplitudes [89]. For a thorough study of symmetries in the context of the SFA, see [80]. Furthermore, although the CQSFA has been successfully used to study rare gases or small molecules [46, 68, 66, 67] employing effective potentials, the extra phase shifts arising from them are not useful to our problem. They have been used, however, to detect the parity of orbitals [68]. Different targets will also introduce additional momentum biases stemming from the bound-state geometries [90, 83, 91].
@@
In the present paper, we have found that the agreement between the results obtained using our theories is fairly good if the solutions used in the CQSFA theory are restricted to only those which we expect to have a counterpart in the SFA theory. This happens because the compared orbits accumulate similar phases in the continuum, although their dynamics can be very different. For example, before the rescattering, the SFA trajectories are one-dimensional for a linearly polarized driving field, while for the CQSFA theory, the corresponding orbits are always two-dimensional. More generally, the SFA trajectories follow the field, which is not the case with their CQSFA counterparts due to the always present Coulomb interaction.
@@ Acknowledgements.
- Schafer et al. [1993] K. J. Schafer, B. Yang, L. F. DiMauro, and K. C. Kulander, Above threshold ionization beyond the high harmonic cutoff, Phys. Rev. Lett. 70, 1599 (1993).
@@
- Spanner et al. [2004] M. Spanner, O. Smirnova, P. B. Corkum, and M. Y. Ivanov, Reading diffraction images in strong-field ionization of diatomic molecules, J. Phys. B 37, L243 (2004).
- Maxwell et al. [2021] A. S. Maxwell, A. Serafini, S. Bose, and C. F. d. M. Faria, Quantum estimation in strong fields: In situ ponderomotive sensing, Phys. Rev. A 103, 043519 (2021).
- Maxwell and Figueira de Morisson Faria [2015] A. S. Maxwell and C. Figueira de Morisson Faria, Quantum interference in time-delayed nonsequential double ionization, Phys. Rev. A 92, 023421 (2015).
- Maxwell and Figueira de Morisson Faria [2016] A. S. Maxwell and C. Figueira de Morisson Faria, Controlling below-threshold nonsequential double ionization via quantum interference, Phys. Rev. Lett. 116, 143001 (2016).
- Werby et al. [2021] N. Werby, A. S. Maxwell, R. Forbes, P. H. Bucksbaum, and C. Figueira de Morisson Faria, Dissecting subcycle interference in photoelectron holography, Phys. Rev. A 104, 013109 (2021).
- Werby et al. [2022] N. Werby, A. S. Maxwell, R. Forbes, C. Figueira de Morisson Faria, and P. H. Bucksbaum, Time correlation filtering reveals two-path electron quantum interference in strong-field ionization, arXiv.2205.07166 [physics.atom-ph] (2022).
- Kang et al. [2020] H. Kang, A. S. Maxwell, D. Trabert, X. Lai, S. Eckart, M. Kunitski, M. Schöffler, T. Jahnke, X. Bian, R. Dörner, and C. Figueira de Morisson Faria, Holographic detection of parity in atomic and molecular orbitals, Phys. Rev. A 102, 013109 (2020).
@@
- Maxwell et al. [2022] A. S. Maxwell, L. B. Madsen, and M. Lewenstein, Entanglement of orbital angular momentum in non-sequential double ionization, Nature Communications 13, 4706 (2022).
- Borbély et al. [2019] S. Borbély, A. Tóth, D. G. Arbó, K. Tőkési, and L. Nagy, Photoelectron holography of atomic targets, Phys. Rev. A 99, 013413 (2019).
- Murakami and Zhang [2020] M. Murakami and G.-P. Zhang, Observation of attosecond electron dynamics in the photoelectron momentum distribution of atoms using few-cycle laser pulses, Phys. Rev. A 101, 053439 (2020).
- Yuan et al. [2021] M.-H. Yuan, A. D. Bandrauk, and X.-B. Bian, Exploring recollision of ultrafast electrons from photoelectron momentum distributions using single-cycle near-infrared laser pulses, Phys. Rev. A 103, 013108 (2021).
- Taoutioui and Tőkési [2022] A. Taoutioui and K. Tőkési, Identifying the complexity of the holographic structures in strong field ionization, Scientific Reports 12, 2877 (2022).
- Bergues et al. [2011] B. Bergues, S. Zherebtsov, Y. Deng, X. Gu, I. Znakovskaya, R. Kienberger, F. Krausz, G. Marcus, and M. F. Kling, Sub-cycle electron control in the photoionization of xenon using a few-cycle laser pulse in the mid-infrared, New Journal of Physics 13, 063010 (2011).
- Möller et al. [2014] M. Möller, F. Meyer, A. M. Sayler, G. G. Paulus, M. F. Kling, B. E. Schmidt, W. Becker, and D. B. Milošević, Off-axis low-energy structures in above-threshold ionization, Phys. Rev. A 90, 023412 (2014).
- Arbó et al. [2015] D. G. Arbó, C. Lemell, S. Nagele, N. Camus, L. Fechner, A. Krupp, T. Pfeifer, S. D. López, R. Moshammer, J. Burgdörfer, and et al., Ionization of argon by two-color laser pulses with coherent phase control, Physical Review A 92, 10.1103/physreva.92.023402 (2015).
- Krečinić et al. [2018] F. Krečinić, P. Wopperer, B. Frusteri, F. Brauße, J.-G. Brisset, U. De Giovannini, A. Rubio, A. Rouzée, and M. J. J. Vrakking, Multiple-orbital effects in laser-induced electron diffraction of aligned molecules, Phys. Rev. A 98, 041401 (2018).
- Bray et al. [2021] A. C. Bray, A. S. Maxwell, Y. Kissin, M. Ruberti, M. F. Ciappina, V. Averbukh, and C. Figueira de Morisson Faria, Polarization in strong-field ionization of excited helium, Journal of Physics B: Atomic, Molecular and Optical Physics 54, 194002 (2021).